A 21.64 g sample of a nonreactive metal is placed in a flask containing 12.00 mL of water; the final volume is 13.81 mL. If the length of the edge of the unit cell is 387 pm and the metallic radius is 137 pm, determine the packing arrangement and identify the element.


1
Expert's answer
2022-03-02T03:10:02-0500

Volume of metal "=13.81-12=1.81ml"


Density of metal "=\frac{21.64}{1.81}=11.9558g/mol"